# Actminer: Applying Causality Tracking and Increment Aligning for Graph-based Threat Hunting

### v1

@@ V Evaluation
  • RQ4: How efficient is Actminer compared with the SOTA in terms of runtime overhead?
  • RQ5: How robust is Actminer in benign dataset?
DataSet. Actminer is evaluated using three datasets: DARPA E3 [20], DARPA E4, Simulated Environments. The DarpaE3 dataset is open-source, whereas the DarpaE4 dataset is not publicly available. Theia and FiveDirections are both from DARPA Engagement 3, and Trace is from DARPA Engagement 3 and 4. The data of Theia and Trace was collected from the Linux, while the data of FiveDirections was collected from the Windows 7. As shown in Table V, the duration encompasses both benign activities and attack-related activities within a dataset, wherein only a small portion of the total time frame involves actual attacks. Detailed description of both actual attack behaviors and benign operations can be found in the Table IV.
As show in Table IV, these attacks include the following scenarios: malicious file downloads and execution, information collection and exfiltration, Firefox memory attacks, backdoor extensions, and phishing emails. Furthermore, we employ Kellect [34] to extract benign data from the Windows platform, while we utilize SPADE [38] for the acquisition of benign log data from Linux. All benign data includes various daily system operations, such as network browsing, file operations, etc. We also use the OpTC dataset which cotnains benign activities of 500 Windows hosts over seven days. These benign datasets contain billions of audit records on Windows, Linux, and FreeBSD.
TABLE IV: The detail of the attack and benign datasets.
@@
TABLE V: The summary of the experimental dataset. Column 1 specifies the name of dataset, and Column 2 denotes the corresponding duration. Columns 3 and 4 indicate the number of nodes and edges, respectively. Column 5 represents the number of attack nodes.
@@
Detector for Comparison. To evaluate Actminer, we use a graph alignment-based threat hunting system as benchmark: POIROT. Why are we only comparing Actminer to POIROT? There are already several threat hunting studies [2, 39, 25, 11, 40]. We broadly categorise them into the following two based on the techniques they use: machine learning-based approaches [39, 11, 40] and search-based approaches [2, 25]. While both DeepHunter [39] and ProvG-Searcher [11] conduct model training by constructing positive (attack graphs extracted manually from the provenance graph) and crafted negative samples. MEGR-APT[40] utilizes a graph matching model to compute similarity scores between the query graph’s embedding vector and the embedding vectors of detected subgraphs. However, a fundamental limitation of these coarse-grained methods is their inability to consider the relationships between nodes in the attack chain (e.g., DeepHunter solely considers the relationships between IOCs rather than the comprehensive association information of all attack nodes). In other words, the results obtained from DeepHunter and ProvG-Searcher may not necessarily represent complete attack chains. Furthermore, our initial attempt to re-implement the ProvG-Searcher revealed that the core component responsible for processing both provenance graphs and query graphs is not available as an open-source solution. Hence, we do not compare our work with them. ThreatRaptor [25] uses NLP technology to extract threat behaviour graphs from CTI reports and transforms the graphs into TBQL query language using specific algorithms. Unlike ACTMINER, it stores audit log data in a database, allowing for the retrieval of individual attack behaviors. The single-point matching results obtained from the TBQL query statement do not correspond well to the contextual content of the attack process described in the CTI report. So we do not compare with it.
Here, we compare the performance of Actminer with POIROT [2], which are the most relevant in term of level and methodology for our evaluation. POIROT aligns the query graph with the provenance graph based on node type and name regularization. Due to the unavailability of the query graphs manually extracted by the authors in POIROT, for fairness, our query graph is uniformly extracted by the Extractor [28] in our experimental setting. However, during the extraction process, we encounter graph disconnections, incomplete attack chains, etc. Therefore, we use the state-of-the-art method CRUcialG [41] to assist in obtaining the query graph.
@@
V-A RQ1: How effectively can Actminer detect the attacks especially in terms of false alarms?
@@
Table VI presents the performance of Actminer and POIROT on our evaluation datasets. Actminer consistently surpass POIROT, achieving superior precision, recall values and lower number of FN/FP. In comparison to POIROT, Actminer utilizes entities attributes and action semantic to generate a more generalizable query graph. This provides brief abstracted entity information to taking a graph alignment, subsequently reducing false negative and enhancing precision and recall. As the POIROT paper lacks evaluation on E4 dataset, we execute POIROT on E4 to obtain evaluation results. The findings demonstrate that Actminer significantly outperforms POIROT, as E4 attacks more challenging to detect due to well-blended malicious activity. POIROT relies solely on name and simple types as query graph’s features, which is insufficient. On the other hand, the entities attributes and action semantic employed by Actminer offer more semantic information for each node and less sensitive for nodes with certain types of characteristics, making it harder for attack nodes to conceal themselves.
@@
At first sight, Actminer shows incremental improvement in comparison to POIROT in terms of FN. This is attributed to Actminer considering the correlation between semantics of multi-hop nodes. Table VI shows the performance of Actminer and POIROT on all datasets. Actminer does not miss any malicious nodes, i.e., Actminer’s average false negatives (FNs) are 0, reduced by 61 compared to POIROT. On average, the false postive nodes generated by Actminer ($\sim$ 389 nodes) is 1.91 $\times$ less than POIROT. Actminer demonstrates a notable improvement in precision over POIROT, achieving a 2.96% higher precision score, while also outperforming POIROT in terms of recall with a 1.94% increase.
@@
TABLE VI: Performance of Actminer and POIROT. FN denotes the false negative, which occurs when a genuine attack pattern is incorrectly classified as benign. Conversely, FP represents the false positive, where a benign event or data point is mistakenly identified as an attack one. The notation Prec. denotes precision.
@@
V-B RQ2: How robust is Actminer against adversarial attacks?
When an attack occurs, the attacker’s behavior pattern may be highly similar or even nearly identical to the normal system behavior in a regular environment. From a technical perspective, attackers can mimic normal processes at the API call level or employ code injection techniques to make their behavior patterns indistinguishable from normal processes at the low-level system log. This poses a challenge for provenance-based threat hunting. However, by considering richer contextual semantics, differentiation can still be achieved. Goyal et al. [42] devise three strategies for adversarial detection of anomaly detection systems based on graph-level granularity. Enlightened by their work, we design experiments to assess Actminer’s resilience against adversarial attacks.
To access Actminer’s resilience against adversarial attacks, we perform adversarial mimicry attacks on provenance-based graph alignment threat hunting system. To evaluate Actminer’s anti-attack capability, based on the Darpa datasets as a reference, we modify and add attack steps in the provenance graph, primarily considering two scenarios.
Scenario I: The attacker inserts a large number of invalid attack paths into the actual attack chain, attempting to disrupt and mislead the detection algorithm. Test results show that POIROT suffered severe missed detections in this scenario, while Actminer successfully detected all real attack nodes, resisting the attacker’s disruptive attacks.
Scenario II: The attacker uses normal programs to perform operations similar to attack patterns, attempting to introduce false positives. Tests found that the POIROT exhibits varying degrees of false positives, marking normal processes as attack nodes. Our system, on the other hand, effectively distinguishes the true intent of normal programs through behavior association and intent analysis, avoiding false positives. Based on the above two scenarios, we designed the following three strategies:
@@ V-B RQ2: How robust is Actminer against adversarial attacks?
    Strategy I: Insert additional unrelated benign process read/write file flows between process and file read/write operations. As shown in Figure 5;
@@
    Strategy II: Insert additional benign processes between process and process execution operations. As shown in Figure 6;
@@
    Strategy III: Insert additional processes communicating with sockets between processes and socket communication operations. As shown in Figure 7.
@@
The experimental results, as illustrated in Table VII, depict the percentage of nodes and edges incrementally added to the attack graph using the three aforementioned strategies on the Darpa 4 dataset across all attacks. Both nodes and edges are incrementally added at a constant rate.
@@
In terms of recall rate, our system consistently maintains a high average ratio as the proportion of added nodes increases, with a decreasing trend approaching approximately 1%. This characteristic is determined by the design of our system. In contrast, POIROT exhibits a diminishing recall rate as the proportion increases, with the rate of decline accelerating with higher proportions added. This can be observed from the AVG of R., where the descent shifts from an initial 1.5% at full scale to around 4% towards the latter end of the scale. Regarding precision, as the proportion increases, both our system and POIROT reach their respective lowest points of 89.33% and 69.75%. Through further analysis, we conclude that our system is capable of detecting evasion attacks based on the equivalent semantic approach we have devised.
@@
Why can’t POIROT defend against adversarial attacks? The attacker inserts a large number of irrelevant nodes and connecting edges between the real attack steps, increasing the hop count of a single attack stage to more than three hops. This causes POIROT to fail to associate different stages as a single attack event, resulting in missed detection. POIROT decomposes the graph through bounded branches ($\gamma$) and depth ($\beta$), separating graph neighborhoods from each other. The authors of POIROT consider the optimal depth to be three. Therefore, even if suspicious nodes have a common malicious ancestor, if the hop distance of that behavior from the earliest attack node exceeds $\beta$ hops, the anomaly score obtained by that behavior will be lower than the score of the previous malicious behavior, and that behavior may be mistaken as benign. The attacker can also insert a small number of irrelevant nodes between attack stages, keeping the hop number of the attack chain for each stage within POIROT’s detection threshold of three. Although the inserted nodes themselves are harmless, the edges connecting them to the real attack nodes cause POIROT to erroneously judge them as part of the attack, resulting in false negatives. Through our further investigation, we analyze the reasons behind the superior performance of Actminer: POIROT uses the query graph directly for regular matching of node types and their attribute names, and selects appropriate nodes based on path scores, neglecting to consider temporal and causal issues between nodes, resulting in many false alarms.
@@
The robustness of Actminer results from two reasons. First of all, during the construction of suspicious semantic tree, attack intent information would be confirmed. The asserted behaviors like fork/clone will not influence the score of the certain attack path, because the multiple-hops strategy will consider these nodes to contain the same suspicious semantic. Second, Actminer do not raise alarms unless nodes exceeding the attack sequence length are detected. This ensures that our system does not generate excessive false positives when attackers intend to conduct multi-point blasting to affect the hunting system, while also enhancing its resilience against Scenario I.
@@
TABLE VII: The performance comparison between the Actminer and POIROT across four scenarios (i.e., four query graphs) from the E4-Trace dataset, where varying proportions of edges and nodes are added in the provenance graph. Each addition consists of an equal number of three types of nodes/edges. The percentage added represents the percentage of the number of nodes in the provenance graph relative to the attack node.
@@
V-C RQ3: How important are the components we design for assisting threat hunting?
To demonstrate the impact of our system components, particularly the query graph processing (QGP) and equivalent semantic transfer (EST) components on system performance and efficiency, we conduct ablation experiments on them separately. The QGP step is responsible for generating the required query graphs, and facilitating the subsequent operations of our system. On the other hand, the EST step plays a crucial role in threat hunting module.
1) Efficacy of Equivalent Semantic Transfer. We test the performance of Actminer without EST using the same settings in the RQ1. As demonstrated in Table VI, without the EST component the number of FN enhanced by 1.4%, the number of false positives is 83.96% of POIROT’s, while it represented an increase compared to Actminer. It is attributed to the following reasons: The process of equivalent semantic transfer also inherently involves the causal logic and temporal relationships of the attack to some extent. Therefore, removing this component may introduce a small portion of false positives.
2) Effectiveness of Query Graph Processing. As part of our study, we directly utilize the query graphs extracted by the Extractor and CRUcialG on the same provenance graphs used in RQ1 for threat hunting. The size of query graph before QGP and after QGP for all situations are presented in Table VIII, along with the results and the corresponding analysis. In the QGP process, we uses the method of attribute abstraction for nodes, so the number of nodes would be reduced. Our analysis explores the impact of incorporating Query Graph Processing (QGP) on the system’s accuracy and timeliness. We use the Trace dataset from E3 to study the effect of QGP. The result, presented in Figure 8, shows a significant reduction in time consumption when considering QGP. This improvement results from QGP further merging the extracted nodes. Ignoring query graph processing makes efficiently hunting these nodes challenging. However, incorporating QGP allows the system to identify these nodes more effectively, leading to improved threat hunting performance.

### v2

@@ 6 Evaluation
  • RQ4: How efficient is Actminer compared with SOTA in terms of runtime overhead?
  • RQ5: How robust is Actminer in a benign dataset?
DataSet. Actminer is evaluated using three datasets: DARPAE3 DARPA3 , DARPAE4, and Simulated Environments. DARPAE3 is an open-source dataset, whereas the DARPAE4 dataset is not publicly available. Theia and FiveDirections are both from DARPA Engagement 3, while Trace is from DARPA Engagements 3 and 4. The data of Theia and Trace was collected from Linux, while that of FiveDirections was collected from Windows 7. As presented in Table 5, the duration encompasses both benign activities and attack-related activities within a dataset, wherein only a small portion of the total time frame involves actual attacks. Detailed description of both actual attack behaviors and benign operations is presented in Table 4.
As presented in Table 4, these attacks include the following scenarios: malicious file downloads and execution, information collection and exfiltration, Firefox memory attacks, backdoor extensions, and phishing emails. Furthermore, we employ Kellect chen2022kellect to extract benign data from the Windows platform, while we use SPADE gehani2012spade to acquire benign log data from Linux. All benign data includes various daily system operations, such as network browsing and file operations. We also use the OpTC dataset which contains benign activities of 500 Windows hosts over 7 days. These benign datasets contain billions of audit records on Windows, Linux, and FreeBSD.
Table 4: The detail of the attack and benign datasets.
@@
Table 5: The summary of the experimental dataset. Column 1 specifies the name of dataset, and Column 2 denotes the corresponding duration. Columns 3 and 4 indicate the number of nodes and edges, respectively. Column 5 represents the number of attack nodes.
@@
Detector for Comparison. We first evaluate the performance of Actminer by comparing it with the following state-of-the-art threat hunting methods, including both graph alignment–based and machine learning–based approaches:
  • MEGR-APT MEGR-APT : It proposes a graph neural network–based attack representation learning approach to effectively detect suspicious subgraphs that match attack query patterns. It extracts comprehensive embeddings for both query and provenance graphs by taking node types as initial features and then utilizes a neural tensor network for graph alignment.
  • ProvG-Searcher altinisik2023provg : It introduces a graph simplification strategy that preserves diverse behaviors in provenance graphs, enhancing learning and search capabilities. It leverages order embeddings for efficient evaluation of subgraph relationships in the embedding space.
  • POIROT milajerdi2019poirot : It advances APT attack detection by modeling threat hunting as an inexact graph pattern matching problem that leverages CTI. Through similarity-based alignment between these graph representations, POIROT enables effective detection of suspicious activities that match known attack patterns, even when the match is not exact.
Since POIROT is not open source, we have attempted to implement the steps outlined in the POIROT paper using the algorithm and used point-level ground truth to evaluate POIROT and Actminer. Based on MORSE hossain2020combating , we conduct manual annotation of the DARPAE4 dataset. For the DARPAE3 dataset, we acquired the GT labels from existing studies: Karios cheng2023kairospracticalintrusiondetection , Flash rehman2024flash , ThreaTrace wang2022threatrace , and REAPr reapr-ground-truth . Due to the unavailability of the query graphs manually extracted by the authors in POIROT, for fairness, our query graph is uniformly extracted by the Extractor satvat2021extractor in our experimental setting. However, during the extraction process, we encounter graph disconnections and incomplete attack chains, among other problems. Therefore, we use the state-of-the-art method CRUcialG cheng2024crucialg to assist in obtaining the query graph.
@@
6.1 RQ1: How effectively can Actminer detect the attacks especially in terms of false alarms?
@@
Table 6 presents the performance of Actminer and POIROT on our evaluation datasets. Actminer consistently surpasses POIROT, achieving superior precision, recall values, and lower number of FN/FP. Compared with POIROT, Actminer utilizes entity attributes and action semantics to generate a more generalizable query graph. This provides brief abstracted entity information to take a graph alignment, subsequently reducing FNs and enhancing precision and recall. As the POIROT paper lacks evaluation on E4 dataset, we execute POIROT on E4 to obtain evaluation results. The findings demonstrate that Actminer significantly outperforms POIROT, as E4 attacks are more challenging to detect due to well-blended malicious activity. POIROT relies solely on name and simple types as the query graph’s features, which is insufficient. In contrast, the entity attributes and action semantics employed by Actminer offer more semantic information for each node and are less sensitive for nodes with certain types of characteristics, making it harder for attack nodes to conceal themselves.
@@
At first sight, Actminer shows incremental improvement compared with POIROT in terms of FN. This is attributed to Actminer considering the correlation between the semantics of multihop nodes. Table 6 summarizes the performance of Actminer and POIROT on all datasets. Actminer does not miss any malicious nodes, i.e., its average FNs are 0, reduced by 61 compared to POIROT. On average, the FP nodes generated by Actminer ($\sim$ 389 nodes) is 1.91 $\times$ less than that of POIROT. Actminer demonstrates a notable improvement in precision over POIROT, achieving a 2.96% higher precision score while also outperforming POIROT in terms of recall with a 1.94% increase.
We reimplemented MEGR-APT and ProvG-Searcher and conducted a comparative evaluation. As presented in Table 7, Actminer consistently outperforms both MEGR-APT MEGR-APT and ProvG-Searcher altinisik2023provg across all evaluated datasets and metrics. It also achieves superior precision scores, ranging from 0.95 to 0.98, while maintaining perfect recall (1.00) in most cases. Its area under the curve values ranges from 0.97 to 0.99, demonstrating robust discriminative performance that surpasses the 0.96–0.98 and 0.97–0.98 ranges achieved by MEGR-APT and ProvG-Searcher, respectively.
MEGR-APT’s consideration of temporal sequences and causal relationships in query graphs, coupled with their graph sampling method for threat hunting, contributes to competitive performance; however, it suffers from significant limitations. The system requires substantial preprocessing overhead—approximately 1 h daily for constructing and loading the traceability graph into the RDF database. This preprocessing bottleneck, combined with lower precision scores (MEGR-APT’s 0.93–0.95 vs. Actminer’s 0.95–0.98), impairs real-time performance and limits practical applicability in time-critical security scenarios. Similarly, while ProvG-Searcher demonstrates consistent performance across datasets, it cannot match the precision advantages offered by Actminer, particularly evident in the E4 Trace dataset, in which our system achieves a precision of 0.98 compared with that of ProvG-Searcher, which is 0.97.
@@
Table 6: Performance of Actminer and POIROT. FN denotes the false negative, which occurs when a genuine attack pattern is incorrectly classified as benign. Conversely, FP represents the false positive, where a benign event or data point is mistakenly identified as an attack one. The notation Prec. denotes precision.
@@
Table 7: Comparison of ActMiner with GNN-based threat hunting systems across different datasets.
DataSet System Precision Recall AUC TPR FPR
E4-Trace ActMiner 0.98 1.00 0.99 1.00 0.02
MEGR-APT 0.95 1.00 0.98 1.00 0.04
ProvG-Searcher 0.97 0.99 0.98 1.00 0.03
E3-Trace ActMiner 0.96 1.00 0.98 1.00 0.04
MEGR-APT 0.95 1.00 0.98 1.00 0.03
ProvG-Searcher 0.95 1.00 0.98 1.00 0.03
E3-Theia ActMiner 0.95 1.00 0.97 1.00 0.06
MEGR-APT 0.93 1.00 0.98 1.00 0.04
ProvG-Searcher 0.94 1.00 0.98 1.00 0.03
6.2 RQ2: How robust is Actminer against adversarial attacks?
When an attack occurs, the attacker’s behavior pattern may be highly similar or even nearly identical to the normal system behavior in a regular environment. From a technical perspective, attackers can mimic normal processes at the API call level or employ code injection techniques to make their behavior patterns indistinguishable from normal processes at the low-level system log. This poses a challenge for provenance-based threat hunting. However, differentiation can still be achieved by considering richer contextual semantics. Goyal et al. goyal2023sometimes devise three strategies for adversarial detection of anomaly detection systems based on graph-level granularity. Enlightened by their work, we design experiments to assess Actminer’s resilience against adversarial attacks.
To assess Actminer’s resilience against adversarial attacks, we perform adversarial mimicry attacks on provenance-based graph alignment threat hunting system. To evaluate its antiattack capability, using DARPA datasets as a reference, we modify and add attack steps in the provenance graph, primarily considering two scenarios.
Scenario I: The attacker inserts a large number of invalid attack paths into the actual attack chain, attempting to disrupt and mislead the detection algorithm. Test results show that POIROT exhibited severe missed detections in this scenario, while Actminer successfully detected all real attack nodes, resisting the attacker’s disruptive attacks.
Scenario II: The attacker uses normal programs to perform operations similar to attack patterns, attempting to introduce FPs. Test results showed that POIROT exhibits varying degrees of FPs, marking normal processes as attack nodes. In contrast, Actminer effectively distinguishes the true intent of normal programs through behavior association and intent analysis, avoiding FPs. Based on the above two scenarios, we designed the following three strategies:
@@ 6.2 RQ2: How robust is Actminer against adversarial attacks?
    Strategy I: Insert additional unrelated benign process read/write file flows between process and file read/write operations (Figure 7).
@@
    Strategy II: Insert additional benign processes between process and process execution operations (Figure 8).
@@
    Strategy III: Insert additional processes communicating with sockets between processes and socket communication operations (Figure 9).
@@
The experimental results, as presented in Table 8, depict the percentage of nodes and edges incrementally added to the attack graph using the three aforementioned strategies on the DARPA4 dataset across all attacks. Both nodes and edges are incrementally added at a constant rate.
@@
In terms of recall rate, Actminer consistently maintains a high average ratio as the proportion of added nodes increases, with a decreasing trend approaching approximately 1%. This characteristic is determined by the design of Actminer. In contrast, POIROT exhibits a diminishing recall rate as the proportion increases, with the rate of decline accelerating as the proportions of added nodes increase. This can be observed from the AVG of recall, where the descent shifts from an initial 1.5% at full scale to around 4% toward the end of the scale. In terms of precision, as the proportion increases, both Actminer and POIROT reach their lowest points of 89.33% and 69.75%, respectively. Through further analysis, we conclude that Actminer can detect evasion attacks based on the equivalent semantic approach we have devised.
@@
Why can POIROT not defend against adversarial attacks? The attacker inserts a large number of irrelevant nodes and connecting edges between the real attack steps, increasing the hop count of a single attack stage to more than three hops. This causes POIROT to fail to associate different stages as a single attack event, resulting in missed detection. POIROT decomposes the graph through bounded branches ($\gamma$) and depth ($\beta$), separating graph neighborhoods from each other. The authors of POIROT consider the optimal depth to be three. Therefore, even if suspicious nodes have a common malicious ancestor, if the hop distance of that behavior from the earliest attack node exceeds $\beta$ hops, the anomaly score obtained by that behavior will be lower than the score of the previous malicious behavior, and that behavior may be mistaken as benign. The attacker can also insert a small number of irrelevant nodes between attack stages, keeping the hop number of the attack chain for each stage within the detection threshold of POIROT to three. Although the inserted nodes themselves are harmless, the edges connecting them to the real attack nodes cause POIROT to erroneously judge them as part of the attack, resulting in FNs.
Through further investigation, we analyze the reasons behind the superior performance of Actminer: POIROT uses the query graph directly for regular matching of node types and their attribute names and selects appropriate nodes based on path scores, neglecting to consider temporal and causal issues between nodes, resulting in many false alarms.
@@
The robustness of Actminer results from two reasons. First, attack intent information would be confirmed during the construction of suspicious semantic tree. The asserted behaviors such as fork/clone will not influence the score of the certain attack path, because the multiple hops strategy will consider these nodes to contain the same suspicious semantic. Second, Actminer does not raise alarms unless nodes exceeding the attack sequence length are detected. This ensures that Actminer does not generate excessive FPs when attackers intend to conduct multipoint blasting to affect the hunting system while also enhancing its resilience against Scenario I.
@@
Table 8: The performance comparison between the Actminer and POIROT across four scenarios (i.e., four query graphs) from the E4-Trace dataset, where varying proportions of edges and nodes are added in the provenance graph. Each addition consists of an equal number of three types of nodes/edges. The percentage added represents the percentage of the number of nodes in the provenance graph relative to the attack node.
@@
6.3 RQ3: How important are the components we design for assisting threat hunting?
To demonstrate the impact of our system components, particularly the query graph processing (QGP) and EST components on system performance and efficiency, we conduct ablation experiments on them separately. The QGP step is responsible for generating the required query graphs and facilitating the subsequent operations of our system. In comparison, the EST step plays a crucial role in threat hunting module.
1) Efficacy of EST. We test the performance of Actminer without EST using the same settings in RQ1. As presented in Table 6, without the EST component, Actminer achieved an improved FNs by 1.4% and exhibited FP that is 83.96% to that of POIROT’s, while it represented an increase compared to Actminer.
It is attributed to the following reasons: The EST process also inherently involves the causal logic and temporal relationships of the attack to some extent. Therefore, removing this component may introduce a small portion of FPs.
2) Effectiveness of QGP. As part of our study, we directly utilize the query graphs extracted by the Extractor and CRUcialG on the same provenance graphs used in RQ1 for threat hunting. The size of the query graph before and after QGP for all situations is presented in Table 9, along with the results and the corresponding analysis. In the QGP process, we use the method of attribute abstraction for nodes, resulting in a reduced number of nodes. Our analysis explores the impact of incorporating QGP on the accuracy and timeliness of the system. We use the Trace dataset from E3 to study the effect of QGP. The result, as depicted in Figure 10, shows a significant reduction in time consumption when considering QGP. This improvement results from QGP further merging the extracted nodes. Ignoring QGP makes it challenging to efficiently hunt these nodes. However, incorporating QGP allows the system to identify these nodes more effectively, leading to improved threat hunting performance.